NuFloor Manufacturing produces floor tiles. The managers at NuFloor are trying to develop budgets for the upcoming quarter. The following data have been gathered for a particular tile design.

  

Projected unit sales 1,890 cases
Selling price per case $ 240
Inventory at the beginning of the quarter 150 cases
Target inventory at the end of the quarter 100 cases
Direct labor hours needed to produce one case 2 hours
Direct labor wages $ 10 per hour
Direct material cost per case $ 8
Variable manufacturing overhead cost per case $ 6
Fixed overhead costs for the upcoming quarter $ 220,000

What is the projected ending inventory value for this particular tile design?

Ending Inventory Value = $
Answer #1

Projected Production During the Quarter = Estimated Sales for the Quarter + Ending Inventory of Finished Goods - Beginning Inventory of Finished Goods

Projected Production during the Quarter = 1,890 + 100 - 150

Projected Production during the Quarter = 1,840 Units

Using Absorption Costing

Treating Fixed Overhead Costs as part of Product Costs

Particulars Amount
Direct Materials Cost per Case 8
Direct Labour Cost per Case 20
Variable Overhead Cost per Case 6
Fixed Overhead Cost per Case 119.57
Unit Product Cost per Case 153.57

Fixed Overhead Cost per Case = 220,000 / 1,840 Units Produced = $ 119.57 per Unit

Direct Labour Cost = 2 Hours per Case * $ 10 per Hour = $ 20

Value of Ending Inventory = 100 Units * 153.57 Per Unit = $ 15,357

Using Variable Costing

Treating Fixed Overhead Cost as Period Cost

Partculars Amount
Direct Materials Cost per Unit 8
Direct Labour Cost per Unit 20
Variable Manufacturing Overhead per Unit 6
Unit Product Cost 34

Value of Ending Inventory as per Variable Costing = 100 Units * $ 34 per Unit = $ 3,400

Additional Notes

I have solved the Question using both Variable and Absorption Costing as there was nothing specifically mentioned so please consider the one which is more relevant as per requirement.